OpenStudy (anonymous):

Find the percent of the number. 113% of 140 A. 15,820 B. 1.582 C. 1582 D. 158.2

OpenStudy (anonymous):

113% of 140= 140*1.13 =158.2

sammixboo (sammixboo):

Please do not give out the direct answer =) Guide the user to the answer, so the user can learn how to do the problem for future similar problems =)

OpenStudy (anonymous):

she did guide me

OpenStudy (anonymous):

32 is 80% of what number? A. 0.4 B. 4 C. 25.6 D. 40

OpenStudy (anonymous):

My apologies @sammixboo. @fgmghhbjmb, please disregard the last part of my answer. If you have questions as to how to solve for percentages, please let me know.

sammixboo (sammixboo):

Well, you can use the formula \(\rm\color{blueviolet}{\dfrac{part}{whole}=\dfrac{percent}{100}}\) We now that our 32 is our part and 80% is our percent, so plug in the values \(\rm\color{blueviolet}{\dfrac{32}{whole}=\dfrac{80}{100}}\) We are solving for the whole. Do you know how to get that?